OpenStudy (anonymous):

If h(x) = −4x − 10, find h(−5)

OpenStudy (anonymous):

-4(-5)-10=20-10=10

OpenStudy (anonymous):

h(-5) = -4(-5) - 10 h(-5) = 20 - 10 h(-5) = 10